# Distributional Limits for Eigenvalues of Graphon Kernel Matrices

### v1

@@ Abstract
We study fluctuations of individual eigenvalues of kernel matrices arising from dense graphon-based random graphs. Under minimal integrability and boundedness assumptions on the graphon, we establish distributional limits for simple, well-separated eigenvalues of the associated integral operator. We show that a sharp dichotomy governs the asymptotic behavior. In the non-degenerate regime, the properly normalized empirical eigenvalue satisfies a central limit theorem with an explicit variance, whereas in the degenerate regime the leading fluctuations vanish and the centered eigenvalue converges to an explicit weighted chi-square limit determined by the operator spectrum.
Our analysis requires no smoothness or Lipschitz-type assumptions on the graphon. While earlier work under comparable integrability conditions established operator convergence and eigenspace consistency, the present results characterize the full fluctuation behavior of individual eigenvalues, thereby extending eigenvalue fluctuation theory beyond regimes accessible through operator convergence alone.
Spectral perturbation,
keywords:
[class=MSC]
keywords:
and
@@ 1 Introduction
Spectral methods are among the most widely used tools for extracting structure from network data (von Luxburg, 2007; Ng et al., 2002). Yet despite their ubiquity in community detection (Rohe et al., 2011; Lei and Rinaldo, 2015; Zhao et al., 2012), clustering (Sarkar and Bickel, 2015), and goodness-of-fit procedures (Bickel et al., 2011; Gao et al., 2017), the uncertainty quantification of spectral summaries remains underdeveloped. Even in dense latent-variable models, where eigenvalues are consistent and computationally stable (Koltchinskii and Giné, 2000; Zhu, 2020), it is still not well understood when an individual empirical eigenvalue exhibits Gaussian fluctuations and when it displays fundamentally non-Gaussian behavior. While statistical inference for spectral clustering has been explored in various contexts (Athreya et al., 2017; Tang et al., 2021), including confidence intervals for clustering assignments and goodness-of-fit tests for network models (Levin et al., 2017; Lyzinski et al., 2017), distributional theory for individual eigenvalues under minimal regularity assumptions remains comparatively underdeveloped.
This paper provides a distributional theory for individual eigenvalues in dense graphon-based random graphs. We show that a sharp dichotomy governs eigenvalue fluctuations: depending on the degeneracy structure induced by the population eigenfunction, the leading fluctuations are either asymptotically Gaussian or converge to an explicit weighted chi-square law. This dichotomy was first established by Chatterjee and Huang (2025) under Lipschitz regularity of the graphon; the present work extends their result to the minimal setting of $L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$ kernels, thereby covering discontinuous models including stochastic block models with unequal block sizes. A notable consequence is that, at the $\sqrt{n}$ scale, the dominant randomness comes from sampling the latent positions, while Bernoulli edge noise is asymptotically negligible in non-degeneracy regime (Lei and Rinaldo, 2015; Abbe, 2017; Chin et al., 2015). This separation clarifies the sources of uncertainty in spectral procedures and enables principled inference for eigenvalue-based methods in network analysis.
In the graphon model developed by Lovász and Szegedy (2006); Lovász (2012); Borgs et al. (2008, 2012), each vertex $i$ is associated with a latent variable $U_{i}\in[0,1]$, typically sampled i.i.d. from $\mathrm{Unif}[0,1]$, and edges are drawn independently with $\mathrm{Bernoulli}\bigl(W(U_{i},U_{j})\bigr),i<j,$ for a symmetric measurable kernel $W:[0,1]^{2}\to[0,1]$. This construction induces a random kernel matrix $(\mathbf{K}_{n})_{ij}:=W(U_{i},U_{j})\mathbbm{1}\{i\neq j\}$, and includes canonical models such as stochastic block models Holland et al. (1983); Abbe (2017); Abbe and Sandon (2018) and random dot product graphs Athreya et al. (2017); Levin et al. (2017). Conditionally on the latent positions, $\mathbb{E}[\mathbf{A}_{n}\mid U_{1},\dots,U_{n}]=\mathbf{K}_{n},$ so $\mathbf{K}_{n}$ captures the latent mean geometry while $\mathbf{A}_{n}-\mathbf{K}_{n}$ is pure Bernoulli edge noise. We ask which source governs the leading fluctuations of individual eigenvalues.
From a theoretical perspective, distributional limits for individual eigenvalues provide a refinement of classical spectral convergence results for random kernel matrices and graph Laplacians, which are typically formulated in almost-sure or $L^{2}$ terms Koltchinskii and Giné (2000); Karoui (2010); Braun (2006). Foundational work by Koltchinskii and Giné Koltchinskii and Giné (2000) established operator convergence and eigenspace consistency, together with central limit theorems for certain linear spectral functionals, but sharp fluctuation theory for individual eigenvalues requires additional analysis beyond operator-level convergence. Recently, Chatterjee and Huang (2025) derived a precise eigenvalue fluctuation dichotomy for graphon-based kernel matrices under Lipschitz regularity assumptions on the kernel.
From an applied standpoint, eigenvalue fluctuation results underpin confidence intervals and hypothesis tests for community structure, graph comparison, and goodness-of-fit in network models Bickel et al. (2011); Zhao et al. (2012); Gao et al. (2017). In the graphon setting, however, existing results either rely on strong smoothness conditions or focus primarily on spectral convergence and clustering consistency rather than explicit distributional limits under minimal assumptions Lovász and Szegedy (2006); Wolfe and Olhede (2013); Lei and Rinaldo (2015). As a result, eigenvalue fluctuation theory for graphon-based kernel matrices under very mild regularity conditions remains comparatively underdeveloped.
Our contributions are: (i) a unified fluctuation theory for individual eigenvalues under minimal integrability conditions, (ii) an explicit characterization of the Gaussian versus weighted chi-square regimes via degeneracy, and (iii) a demonstration that $\sqrt{n}$-scale eigenvalue uncertainty is driven by latent-position sampling rather than edge noise.
Our proof strategy combines three main ideas. First, a second-order Rayleigh–Schrödinger expansion for simple eigenvalues, which isolates the leading linear term and accounts for resolvent corrections. Second, concentration of operator norms for bounded kernel matrices, controlling deviations of $\mathbf{K}_{n}$ (and $\mathbf{A}_{n}$ when needed) from their conditional means. Third, Hoeffding-type decompositions of the resulting symmetric statistics, separating the linear, sampling-driven contribution from higher-order degenerate terms. This synthesis clarifies the mechanism behind the Gaussian versus weighted chi-square regimes.
@@
The paper is organized as follows. Section 2 introduces the graphon model, the kernel matrix $\mathbf{K}_{n}$, and the associated integral operator, and fixes notation and normalization. Section 3 states the main fluctuation theorem and outlines the Rayleigh-Schrödinger approach, including implications for the adjacency matrix $\mathbf{A}_{n}$. Section 4 presents non-Lipschitz examples satisfying our assumptions, and Section 5 discusses consequences for spectral methods and directions for future work. Technical proofs, including concentration bounds, Hoeffding decompositions, and resolvent expansions, are deferred to Section 6.
@@ 3 Main Results
Theorem 3.3 (Unified eigenvalue fluctuation dichotomy).
Under Assumption 3.1, let $\lambda_{r}$ denote the $r$th eigenvalue of the population integral operator $T_{W}$. Then:
@@
Remark 3.4 (Origin of the degenerate coefficients).
@@ 3.1 Proof strategy
We outline the main steps in the proof of Theorem 3.3. The analysis is carried out at the matrix level using second-order perturbation theory for simple eigenvalues, combined with Hoeffding decompositions of U-statistics.
@@
By Lemma A.0.1, the $r$th eigenvector of $\mathbb{E}[\widetilde{\mathbf{K}}_{n}]$ is well approximated by the normalized discretization $\mathbf{u}_{r}$, and the corresponding eigengap is bounded below by $\gamma_{r}$. Applying the Rayleigh-Schrödinger expansion for simple eigenvalues yields
@@ 3.2 Implications for the adjacency matrix
Corollary 3.5 (Adjacency eigenvalue fluctuations in the non-degenerate regime).
Let $\mathbf{A}_{n}$ be the adjacency matrix with $\mathbb{E}[\mathbf{A}_{n}\mid U_{1},\dots,U_{n}]=\mathbf{K}_{n}$. Assume Assumption 3.1 and suppose $\sigma_{r}^{2}:=\operatorname{Var}(\varphi_{r}(U)^{2})>0$. Then
@@
The first display follows from the eigenvalue perturbation bound for inhomogeneous Bernoulli graphs; see Theorem 2 in Aalipur and Qin (2026). The second display follows by Slutsky’s theorem, combining the first display with the non-degenerate conclusion of Theorem 3.3 for $\lambda_{r}(\mathbf{K}_{n}/(n-1))$. ∎
Remark 3.6.
If $\varphi_{r}^{2}\equiv 1$ almost everywhere, then the limit in Theorem 3.3 is non-Gaussian, and the same non-Gaussian limit carries over from $\mathbf{K}_{n}$ to $\mathbf{A}_{n}$ under the same edge-noise perturbation bound.
@@ 4 Examples
This section presents several examples of graphons that satisfy Assumption 3.1 and therefore fall within the scope of Theorem 3.3, while lying outside the Lipschitz framework considered in earlier work. These examples illustrate that $\sqrt{n}$-scale eigenvalue fluctuations depend primarily on spectral and integrability properties rather than smoothness, and they clarify the distinction between the non-degenerate and degenerate regimes.
@@ 4.1 Brownian square-root kernel
Nevertheless, $W\in L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$ with $\|W\|_{\infty}\leq 2$. Under Assumption 3.1, Theorem 3.3 applies and yields $\sqrt{n}(\lambda_{r}(\mathbf{K}_{n})/(n-1)-\lambda_{r})\xrightarrow{d}\mathcal{N}(0,\lambda_{r}^{2}\sigma_{r}^{2})$ whenever $\sigma_{r}^{2}>0$, showing coverage beyond Lipschitz kernels.
@@ 4.2 Stochastic block model with unequal block sizes
The kernel satisfies Assumption 3.1. It belongs to $L^{2}([0,1]^{2})$ and $L^{\infty}([0,1]^{2})$, its spectrum is finite dimensional and determined by the block structure, and its eigenfunctions are bounded and piecewise constant. Theorem 3.3 therefore yields
@@ 4.3 Symmetric stochastic block model
By symmetry, the second eigenfunction satisfies $\varphi_{2}(x)\in\{-1,1\}$ almost everywhere, so that $\varphi_{2}^{2}\equiv 1$ and $\operatorname{Var}(\varphi_{2}(U)^{2})=0$. This model therefore lies in the degenerate regime of Theorem 3.3. The eigenvalue fluctuation is non-Gaussian and satisfies
@@
with convergence in $L^{2}$ under the Hilbert-Schmidt condition. Despite its structural similarity to the unequal-block model, symmetry induces a fundamentally different limiting distribution, illustrating the dichotomy underlying Theorem 3.3.
@@ 4.4 $\alpha$-Hölder kernels
Consider the power-law kernel $W(x,y)=(xy)^{\alpha}$ for $\alpha\in(0,1)$. The kernel satisfies Assumption 3.1. It belongs to $L^{2}([0,1]^{2})$ since
@@
cannot hold uniformly on $[0,1]$. The leading eigenfunction $\varphi_{1}$ is strictly positive and non-constant, so that $\operatorname{Var}(\varphi_{1}(U)^{2})>0$. Theorem 3.3 therefore applies in the non-degenerate regime, yielding a Gaussian central limit theorem for $\lambda_{1}(\mathbf{K}_{n})/(n-1)$.
@@ 5 Discussion
This work develops a distributional theory for individual eigenvalues of dense graphon-based random graphs under minimal $L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$ assumptions. Our main result extends the eigenvalue fluctuation dichotomy established by Chatterjee and Huang (2025) beyond Lipschitz kernels and shows that smoothness, while technically convenient in prior analyses, is not intrinsic to the limiting behavior described in Theorem 3.3. The key probabilistic mechanism is instead the degeneracy structure induced by the population eigenfunction and the zero-diagonal kernel-matrix normalization.
A central conceptual takeaway is that, at the $\sqrt{n}$ scale, eigenvalue fluctuations are driven primarily by sampling variability in the latent positions rather than by Bernoulli edge noise. In the non-degenerate regime, where $\sigma_{r}^{2}=\operatorname{Var}(\varphi_{r}(U)^{2})>0$, this leads to a Gaussian limit with variance $\lambda_{r}^{2}\sigma_{r}^{2}$, as in Theorem 3.3. In the degenerate regime, where $\varphi_{r}^{2}\equiv 1$ almost everywhere, the $\sqrt{n}$-scale fluctuations vanish and the centered eigenvalue converges to an explicit weighted chi-square series determined by the spectrum of $T_{W}$. In stochastic block models, this dichotomy cleanly separates unequal-block settings (typically non-degenerate) from symmetric equal-block settings (degenerate), indicating that the non-Gaussian limit reflects an underlying structural symmetry that is invisible to first-order fluctuations.
From an inferential perspective, the results clarify the sources of uncertainty in spectral procedures for network analysis. In the non-degenerate regime, the asymptotic variance of $\lambda_{r}(\mathbf{K}_{n})/(n-1)$ is $\lambda_{r}^{2}\sigma_{r}^{2}/n$, suggesting plug-in estimation strategies for confidence intervals and hypothesis tests for eigenvalue-based methods in latent-position and block models. Moreover, Corollary 3.5 shows that, after normalization, Bernoulli edge noise is asymptotically negligible at the $\sqrt{n}$ scale, so the same distributional limits apply to the observed adjacency matrix $\mathbf{A}_{n}$. The degenerate regime offers complementary diagnostic value: detecting $\varphi_{r}^{2}\equiv 1$ corresponds to identifying symmetry in the generating structure and can inform model checking and selection.

### v2

@@ Abstract
We study the fluctuation behavior of individual eigenvalues of kernel matrices arising from dense graphon-based random graphs. Under minimal integrability and boundedness assumptions on the graphon, we establish distributional limits for simple, well-separated eigenvalues of the associated integral operator. A sharp probabilistic dichotomy emerges: in the non-degenerate regime, the properly normalized empirical eigenvalue satisfies a central limit theorem with an explicit variance, whereas in the degenerate regime the leading stochastic term vanishes and the centered eigenvalue converges to a weighted chi-square law determined by the operator spectrum.
The analysis requires no smoothness or Lipschitz conditions on the kernel. Prior work under comparable assumptions established only operator convergence and eigenspace consistency; the present results characterize the full distributional behavior of individual eigenvalues, extending fluctuation theory beyond the reach of classical operator-level arguments. The proofs combine second-order perturbation expansions, concentration bounds for kernel matrices, and Hoeffding decompositions for symmetric statistics, revealing that at the $\sqrt{n}$ scale the dominant randomness arises from latent-position sampling rather than Bernoulli edge noise.
Keywords: Graphon models; eigenvalue distributions; kernel matrices; U-statistics; spectral perturbation.
MSC2020: Primary 60B20; Secondary 60F05, 62G20, 05C80.
@@ 1 Introduction
Spectral methods play a central role in the analysis of random graphs and network data Lei and Rinaldo (2015); Rohe et al. (2011). Their success in community detection Abbe (2017); Lei and Rinaldo (2015); Zhao et al. (2012), clustering von Luxburg (2007), and goodness-of-fit procedures Bickel et al. (2011); Gao et al. (2015) has motivated a substantial body of work on the asymptotic behavior of eigenvalues and eigenvectors in latent-variable models. Classical results establish almost-sure or $L^{2}$ convergence of empirical kernel operators Koltchinskii and Giné (2000); El Karoui (2008); Braun (2006), together with consistency of eigenspaces and certain linear spectral statistics. However, the distributional behavior of individual eigenvalues, a question of central importance for uncertainty quantification remains comparatively underdeveloped.
This paper develops a distributional theory for simple eigenvalues of kernel matrices arising from dense graphon-based random graphs. Working in the graphon framework of Lovász and Szegedy Lovász and Szegedy (2006), Lovász Lovász (2012), and Borgs et al. Borgs et al. (2008, 2014), we show that the fluctuations of each empirical eigenvalue exhibit a sharp probabilistic dichotomy. Depending on the degeneracy structure of the population eigenfunction, the leading fluctuations are either asymptotically Gaussian or converge to an explicit weighted chi-square law. A version of this dichotomy was first obtained by Chatterjee and Huang Chatterjee and Huang (2021) under Lipschitz regularity of the graphon. Our contribution is to extend this phenomenon to the minimal integrability setting $W\in L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$, thereby covering discontinuous, non-smooth, and piecewise-constant models such as stochastic block models with unequal block sizes. The results demonstrate that smoothness assumptions, while technically convenient, are not intrinsic to the limiting behavior.
From a probabilistic perspective, the fluctuation behavior is driven by the interaction between two sources of randomness: latent-position sampling and Bernoulli edge noise. A central finding of this work is that, at the $\sqrt{n}$ scale, the dominant contribution arises from the latent variables, while edge noise is asymptotically negligible. This separation, implicit in earlier work on spectral convergence Koltchinskii and Giné (2000); Zhu (2020) and latent-position models Lei and Rinaldo (2015); Abbe (2017); Athreya et al. (2017), becomes explicit in our analysis and clarifies the sources of uncertainty in spectral procedures. In particular, the zero diagonal of the kernel matrix-an intrinsic structural feature of graphon models-plays a decisive role in the degenerate regime, where the linear Hoeffding projection vanishes and the limiting distribution becomes a weighted chi-square series.
Our proof combines three probabilistic ingredients. First, a second-order Rayleigh-Schrödinger expansion for simple eigenvalues Kato (1995) reduces the problem to the analysis of a centered Rayleigh quotient. Second, concentration of kernel matrices under boundedness assumptions yields operator-norm control of the empirical perturbation. Third, a Hoeffding decomposition for the resulting symmetric statistic Serfling (1980) isolates the linear and degenerate components, revealing the mechanism behind the Gaussian versus weighted chi-square regimes. This synthesis demonstrates that the fluctuation dichotomy persists well beyond the smooth settings considered in Chatterjee and Huang (2021), and that the limiting distributions are determined by spectral structure and degeneracy rather than regularity of the kernel.
@@
The remainder of the paper is organized as follows. Section 2 introduces the graphon model, the kernel matrix, and the associated integral operator. Section 3 states the main fluctuation theorem and outlines the perturbation-theoretic approach. Section 4 presents non-Lipschitz examples illustrating the scope of the theory. Section 5 discusses implications for spectral methods and directions for future work. Technical proofs are given in Section 3.1.
@@ 3 Main Results
Theorem 3 (Unified eigenvalue fluctuation dichotomy).
Under Assumption 1, let $\lambda_{r}$ denote the $r$th eigenvalue of the population integral operator $T_{W}$. Then:
@@
Remark 4 (Origin of the degenerate coefficients).
@@ 3.1 Proof strategy
We outline the main steps in the proof of Theorem 3. The analysis is carried out at the matrix level using second-order perturbation theory for simple eigenvalues, combined with Hoeffding decompositions of U-statistics.
@@
By Lemma 7, the $r$th eigenvector of $\mathbb{E}[\widetilde{\mathbf{K}}_{n}]$ is well approximated by the normalized discretization $\mathbf{u}_{r}$, and the corresponding eigengap is bounded below by $\gamma_{r}$. Applying the Rayleigh-Schrödinger expansion for simple eigenvalues yields
@@ 3.2 Implications for the adjacency matrix
Corollary 5 (Adjacency eigenvalue fluctuations in the non-degenerate regime).
Let $\mathbf{A}_{n}$ be the adjacency matrix with $\mathbb{E}[\mathbf{A}_{n}\mid U_{1},\dots,U_{n}]=\mathbf{K}_{n}$. Assume Assumption 1 and suppose $\sigma_{r}^{2}:=\operatorname{Var}(\varphi_{r}(U)^{2})>0$. Then
@@
The first display follows from the eigenvalue perturbation bound for inhomogeneous Bernoulli graphs; see Theorem 2 in Aalipur and Qin (2026). The second display follows by Slutsky’s theorem, combining the first display with the non-degenerate conclusion of Theorem 3 for $\lambda_{r}(\mathbf{K}_{n}/(n-1))$. ∎
Remark 6.
If $\varphi_{r}^{2}\equiv 1$ almost everywhere, then the limit in Theorem 3 is non-Gaussian, and the same non-Gaussian limit carries over from $\mathbf{K}_{n}$ to $\mathbf{A}_{n}$ under the same edge-noise perturbation bound.
@@ 4 Examples
This section presents several examples of graphons that satisfy Assumption 1 and therefore fall within the scope of Theorem 3, while lying outside the Lipschitz framework considered in earlier work. These examples illustrate that $\sqrt{n}$-scale eigenvalue fluctuations depend primarily on spectral and integrability properties rather than smoothness, and they clarify the distinction between the non-degenerate and degenerate regimes.
@@ 4.1 Brownian square-root kernel
Nevertheless, $W\in L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$ with $\|W\|_{\infty}\leq 2$. Under Assumption 1, Theorem 3 applies and yields $\sqrt{n}(\lambda_{r}(\mathbf{K}_{n})/(n-1)-\lambda_{r})\xrightarrow{d}\mathcal{N}(0,\lambda_{r}^{2}\sigma_{r}^{2})$ whenever $\sigma_{r}^{2}>0$, showing coverage beyond Lipschitz kernels.
@@ 4.2 Stochastic block model with unequal block sizes
The kernel satisfies Assumption 1. It belongs to $L^{2}([0,1]^{2})$ and $L^{\infty}([0,1]^{2})$, its spectrum is finite dimensional and determined by the block structure, and its eigenfunctions are bounded and piecewise constant. Theorem 3 therefore yields
@@ 4.3 Symmetric stochastic block model
By symmetry, the second eigenfunction satisfies $\varphi_{2}(x)\in\{-1,1\}$ almost everywhere, so that $\varphi_{2}^{2}\equiv 1$ and $\operatorname{Var}(\varphi_{2}(U)^{2})=0$. This model therefore lies in the degenerate regime of Theorem 3. The eigenvalue fluctuation is non-Gaussian and satisfies
@@
with convergence in $L^{2}$ under the Hilbert-Schmidt condition. Despite its structural similarity to the unequal-block model, symmetry induces a fundamentally different limiting distribution, illustrating the dichotomy underlying Theorem 3.
@@ 4.4 $\alpha$-Hölder kernels
Consider the power-law kernel $W(x,y)=(xy)^{\alpha}$ for $\alpha\in(0,1)$. The kernel satisfies Assumption 1. It belongs to $L^{2}([0,1]^{2})$ since
@@
cannot hold uniformly on $[0,1]$. The leading eigenfunction $\varphi_{1}$ is strictly positive and non-constant, so that $\operatorname{Var}(\varphi_{1}(U)^{2})>0$. Theorem 3 therefore applies in the non-degenerate regime, yielding a Gaussian central limit theorem for $\lambda_{1}(\mathbf{K}_{n})/(n-1)$.
@@ 5 Discussion
This work develops a distributional theory for individual eigenvalues of dense graphon-based random graphs under minimal $L^{2}([0,1]^{2})\cap L^{\infty}([0,1]^{2})$ assumptions. Our main result extends the eigenvalue fluctuation dichotomy established by Chatterjee and Huang (2025) beyond Lipschitz kernels and shows that smoothness, while technically convenient in prior analyses, is not intrinsic to the limiting behavior described in Theorem 3. The key probabilistic mechanism is instead the degeneracy structure induced by the population eigenfunction and the zero-diagonal kernel-matrix normalization.
A central conceptual takeaway is that, at the $\sqrt{n}$ scale, eigenvalue fluctuations are driven primarily by sampling variability in the latent positions rather than by Bernoulli edge noise. In the non-degenerate regime, where $\sigma_{r}^{2}=\operatorname{Var}(\varphi_{r}(U)^{2})>0$, this leads to a Gaussian limit with variance $\lambda_{r}^{2}\sigma_{r}^{2}$, as in Theorem 3. In the degenerate regime, where $\varphi_{r}^{2}\equiv 1$ almost everywhere, the $\sqrt{n}$-scale fluctuations vanish and the centered eigenvalue converges to an explicit weighted chi-square series determined by the spectrum of $T_{W}$. In stochastic block models, this dichotomy cleanly separates unequal-block settings (typically non-degenerate) from symmetric equal-block settings (degenerate), indicating that the non-Gaussian limit reflects an underlying structural symmetry that is invisible to first-order fluctuations.
From an inferential perspective, the results clarify the sources of uncertainty in spectral procedures for network analysis. In the non-degenerate regime, the asymptotic variance of $\lambda_{r}(\mathbf{K}_{n})/(n-1)$ is $\lambda_{r}^{2}\sigma_{r}^{2}/n$, suggesting plug-in estimation strategies for confidence intervals and hypothesis tests for eigenvalue-based methods in latent-position and block models. Moreover, Corollary 5 shows that, after normalization, Bernoulli edge noise is asymptotically negligible at the $\sqrt{n}$ scale, so the same distributional limits apply to the observed adjacency matrix $\mathbf{A}_{n}$. The degenerate regime offers complementary diagnostic value: detecting $\varphi_{r}^{2}\equiv 1$ corresponds to identifying symmetry in the generating structure and can inform model checking and selection.